OpenStudy (anonymous):

Choose the equation below that represents the line passing through the point (-2, -3) with a slope of -6. y = -6x - 15 y = -6x - 20 y = -6x + 15 y = -6x + 20

OpenStudy (anonymous):

@welshfella , @Miss.SweetiePie , @Michele_Laino

OpenStudy (welshfella):

the genaral form of this equation is y = mx + b where m = slope and b = where the line cuts the y-axis

OpenStudy (welshfella):

so as you see the slope = -6

OpenStudy (anonymous):

since you know the slope is -6 you can plug that in

OpenStudy (welshfella):

plug in y = -3 and x = -2 into each option and see which one fits

OpenStudy (anonymous):

how would I solve that when plugging them in?

OpenStudy (welshfella):

lets try the second one y = -6x - 20 y = -3, x = -2 so we get left side = -3 right side = -6(-2) - 20 = 12 - 20 = -8 the 2 sides are unequal so it cant be the second equation

OpenStudy (welshfella):

you need the left and right side to be -3

OpenStudy (anonymous):

Okay, so it can't be B right?

OpenStudy (welshfella):

right so try A

OpenStudy (anonymous):

I tried it on paper just like how you did it, and I don't think it is A @welshfella

OpenStudy (anonymous):

I got a

OpenStudy (anonymous):

I'm not sure @welshfella

OpenStudy (welshfella):

a lhs = -3 rhs = -6(-2) - 15 = 12 - 15 = -3 yea A

OpenStudy (anonymous):

Thanks guys!

OpenStudy (welshfella):

yw